# Sample Complexity of Distributionally Robust Average-Reward Reinforcement Learning

### v2

@@ Appendix C Properties of the Bellman Operator: KL-Case
Lemma C.1 (Theorem 1 of Hu and Hong, (2013)).
@@
Our analyses are inspired by the approach in Wang et al., 2024c . To carry out our analysis, we first introduce some notation. As in Wang et al., 2024c , we denote the KL-dual functional under the nominal transition kernel $p_{s,a}$:
@@
Building on Wang et al., 2024c ’s breakthrough in achieving $\delta$-independent bounds through KL-dual analysis, we make two key advances:
@@
From Si et al., (2020), for optimal Lagrange multiplier $\alpha^{*}$, it sufficient to consider $\alpha\in[0,\delta^{-1}\left\|V\right\|_{L^{\infty}(p_{s,a})}]$.
@@ D.3 Anchored DR-AMDP Approach under KL Uncertainty Set
Thus, by Theorem 6 in Wang et al., 2023d , we show $(\xi V_{\mathcal{P}}^{\pi}(s_{0}),V_{\mathcal{P}}^{\pi})$ is a pair of solution to Equation (D.29), where $g=\xi V_{\mathcal{P}}^{\pi}(s_{0})$ is unique. ∎
@@ Appendix F Properties of the Bellman Operator: $f_{k}$-Case
Lemma F.1 (Lemma 1 of Duchi and Namkoong, (2021)).
@@
By Proposition 1 of Duchi and Namkoong, (2021), the dual form of $f(p_{s,a},V,\alpha)$ is concave with respect to $\alpha$, and the supremum is achieved at $\alpha^{*}$ where $\partial_{\alpha}f(p_{s,a},V,\alpha)|_{\alpha=\alpha^{*}}=0$. Since:
@@
$\sup_{\alpha>\left\|V\right\|_{L^{\infty}(p)}}\frac{1}{k^{k}}\left|\frac{\Delta\left[(\alpha-V)^{k^{*}}\right]}{p\left[(\alpha-V)^{k^{*}-1}\right]}\right|\leq\left\|\frac{2\Delta}{p}\right\|_{\infty}\cdot\mathrm{Span}\left(V\right)$ (F.14)
@@
Lemma F.5 (Envelope Theorem, Corollary 3 of Milgrom and Segal, (2002)).
@@ Appendix H Additional Experiment
To further expand on our results in Section 4, we provide additional experiments on baseline comparison and large-scale MDPs. First, we include comparisons with DR RVI Q-learning (Wang et al., 2023e, ) on Hard MDP Instance 5.1 as baseline. Table 2 and Figure 3 shows the error performance for DR RVI Q-learning and the two algorithms for $\mathfrak{p}_{\wedge}=0.9$ and $\delta=0.1$. They demonstrated that the error levels of our two algorithms are comparable and significantly outperform the previous baseline.
@@
Further, to demonstrate the capability of our framework in solving large-scale MDPs, we consider a large-scale MDP with $20$ states and $30$ actions, as in Wang et al., 2023e . The nominal transition distribution is specified by $p_{s,a}\sim\mathcal{N}(1,\sigma_{s,a})$, where $\sigma_{s,a}\sim\textbf{Uniform}[0,100]$, followed by normalization. We then choose the uncertainty size $\delta=0.4$ to introduce stronger perturbations, following the setting in Wang et al., 2023e under the KL-divergence. Note that although $\delta=0.4$ violates Assumption 2, the slope of the linear regression of our results on the logarithmic scale remains very close to 1/2-1/2, further supporting our theoretical guarantees. This observation empirically validates the theoretical results established in our theorems.
Further, to demonstrate the capability of our framework in solving large-scale MDP instances, we consider a large-scale MDP with $20$ states and $30$ actions, as in Wang et al., 2023e . Specifically, Algorithm 2 and Algorithm 3 are evaluated on two distinct large-scale instances, respectively, to verify their effectiveness. The nominal transition distribution is specified by $p_{s,a}\sim\mathcal{N}(1,\sigma_{s,a})$, where $\sigma_{s,a}\sim\textbf{Uniform}[0,100]$, followed by normalization. We then choose the uncertainty size $\delta=0.4$ to introduce stronger perturbations, following the setting in Wang et al., 2023e under the KL-divergence. Note that although $\delta=0.4$ violates Assumption 2, the slope of the linear regression of our results on the logarithmic scale remains very close to 1/2-1/2, further supporting our theoretical guarantees. This observation empirically validates the theoretical results established in our theorems.

### v3

@@ Appendix C Properties of the Bellman Operator: KL-Case
Lemma C.1 (Theorem 1 of Hu and Hong, [15]).
@@
Our analyses are inspired by the approach in Wang et al., 2024c [43]. To carry out our analysis, we first introduce some notation. As in Wang et al., 2024c [43], we denote the KL-dual functional under the nominal transition kernel $p_{s,a}$:
@@
Building on Wang et al., 2024c [43]’s breakthrough in achieving $\delta$-independent bounds through KL-dual analysis, we make two key advances:
@@
From Si et al., [34], for optimal Lagrange multiplier $\alpha^{*}$, it sufficient to consider $\alpha\in[0,\delta^{-1}\left\|V\right\|_{L^{\infty}(p_{s,a})}]$.
@@ D.3 Anchored DR-AMDP Approach under KL Uncertainty Set
Thus, by Theorem 6 in Wang et al., 2023d [44], we show $(\xi V_{\mathcal{P}}^{\pi}(s_{0}),V_{\mathcal{P}}^{\pi})$ is a pair of solution to Equation (D.29), where $g=\xi V_{\mathcal{P}}^{\pi}(s_{0})$ is unique. ∎
@@ Appendix F Properties of the Bellman Operator: $f_{k}$-Case
Lemma F.1 (Lemma 1 of Duchi and Namkoong, [9]).
@@
By Proposition 1 of Duchi and Namkoong, [9], the dual form of $f(p_{s,a},V,\alpha)$ is concave with respect to $\alpha$, and the supremum is achieved at $\alpha^{*}$ where $\partial_{\alpha}f(p_{s,a},V,\alpha)|_{\alpha=\alpha^{*}}=0$. Since:
@@
$\sup_{\alpha>\left\|V\right\|_{L^{\infty}(p)}}\frac{1}{k^{*}}\left|\frac{\Delta\left[(\alpha-V)^{k^{*}}\right]}{p\left[(\alpha-V)^{k^{*}-1}\right]}\right|\leq\left\|\frac{2\Delta}{p}\right\|_{\infty}\cdot\mathrm{Span}\left(V\right)$ (F.14)
@@
Lemma F.5 (Envelope Theorem, Corollary 3 of Milgrom and Segal, [26]).
@@ Appendix H Additional Experiment
To further expand on our results in Section 4, we provide additional experiments on baseline comparison and large-scale MDPs. First, we include comparisons with DR RVI Q-learning [46] on Hard MDP Instance 5.1 as baseline. Table 2 and Figure 3 shows the error performance for DR RVI Q-learning and the two algorithms for $\mathfrak{p}_{\wedge}=0.9$ and $\delta=0.1$. They demonstrated that the error levels of our two algorithms are comparable and significantly outperform the previous baseline.
@@
Further, to demonstrate the capability of our framework in solving large-scale MDPs, we consider a large-scale MDP with $20$ states and $30$ actions, as in Wang et al., 2023e [46]. The nominal transition distribution is specified by $p_{s,a}\sim\mathcal{N}(1,\sigma_{s,a})$, where $\sigma_{s,a}\sim\textbf{Uniform}[0,100]$, followed by normalization. We then choose the uncertainty size $\delta=0.4$ to introduce stronger perturbations, following the setting in Wang et al., 2023e [46] under the KL-divergence. Note that although $\delta=0.4$ violates Assumption 2, the slope of the linear regression of our results on the logarithmic scale remains very close to 1/2-1/2, further supporting our theoretical guarantees. This observation empirically validates the theoretical results established in our theorems.
Further, to demonstrate the capability of our framework in solving large-scale MDP instances, we consider a large-scale MDP with $20$ states and $30$ actions, as in Wang et al., 2023e [46]. Specifically, Algorithm 2 and Algorithm 3 are evaluated on two distinct large-scale instances, respectively, to verify their effectiveness. The nominal transition distribution is specified by $p_{s,a}\sim\mathcal{N}(1,\sigma_{s,a})$, where $\sigma_{s,a}\sim\textbf{Uniform}[0,100]$, followed by normalization. We then choose the uncertainty size $\delta=0.4$ to introduce stronger perturbations, following the setting in Wang et al., 2023e [46] under the KL-divergence. Note that although $\delta=0.4$ violates Assumption 2, the slope of the linear regression of our results on the logarithmic scale remains very close to 1/2-1/2, further supporting our theoretical guarantees. This observation empirically validates the theoretical results established in our theorems.